# On a class of thin obstacle-type problems for the bi-Laplacian operator

### v2

@@ 1. Introduction
$\Delta_{b}w\coloneqq y^{-b}\operatorname{div}(y^{b}\nabla w)=\frac{b}{y}w_{y}+\Delta w,$ (1.2)
@@
$\partial_{y}^{b}w(x,0)\coloneqq\lim_{y\to 0^{+}}y^{b}\partial_{y}w(x,y)$ (1.3)
@@
$f(x,u)\coloneqq\lambda_{-}(u^{-})^{p-1}-\lambda_{+}(u^{+})^{p-1},$ (1.4)
@@ 1.1. Motivation and previous works
$\left\{\begin{array}[]{rll}\Delta_{b}u=&0&\text{ in }B_{1}^{+},\\ \partial_{y}^{b}u=&\lambda_{+}(u^{+})^{p-1}-\lambda_{-}(u^{-})^{p-1}&\text{ on }B_{1}^{\prime}.\end{array}\right$. (1.5)
@@
For more details on the (thick) obstacle problem for the bi-Laplacian, we refer the reader to the works of Frehse (see [27] and [28]), Caffarelli and Friedman (see [10]), and Aleksanyan (see [2]). The regularity of solutions to the thin obstacle problem driven by $\Delta^{2}$ was explored by Schild (see [38] and [37]). More recently, Dipierro, Karakhanyan, and Valdinoci introduced and analyzed a version of the Alt-Caffarelli energy for the bi-harmonic operator (see [18] and [19]).
@@ 1.2. Statement of the main results
    If $b\geq 0$, then $(u,v)\in C^{2,\alpha}(B_{r}^{+})\times C^{0,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,1-b)$;
@@
    If $b<0$, then $(u,v)\in C^{3,\alpha}(B_{r}^{+})\times C^{1,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,-b)$.
@@
    If $b\geq 0$, then $\nabla_{x}^{\prime}v\in C^{0,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,1-b)$;
@@
    If $b<0$, then $\nabla_{x}^{\prime}v\in C^{1,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,-b)$.
@@
Finally, for all $b\in(-1,1)$, there exists $\beta\in(0,1)$ such that $\partial_{y}^{b}v\in C^{0,\beta}(B_{r}^{+})$. In particular, if $b=0$ we have that $(u,v)\in C^{3,\beta}(B_{r}^{+})\times C^{1,\beta}(B_{r}^{+})$ for some $\beta\in(0,1)$.
@@
With Almgren’s frequency formula at our disposal, we follow the pathway outlined in [30]. To be precise, after a preliminary blow-up analysis of the so-called Almgren rescalings of $u$ and $v$ (see ˜6.1), we leverage Weiss- and Monneau-type monotonicity formulas to prove the uniqueness of blow-up solutions. Our findings are summarized in the following theorem, where we use $\mathcal{P}_{\mu}$ to denote the space of $\mu$-homogeneous polynomials that are $b$-harmonic and symmetric with respect to $\{y=0\}$.
@@
for some $\alpha\in(0,1)$ and all $R>0$ as $r\to 0^{+}$. Moreover, the polynomials $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ cannot both vanish identically.
@@ 2.1. Weighted Sobolev Spaces
For a proof of ˜2.3, see Lemma 2.5 in [21].
@@
$\left(\int_{B_{1}^{\prime}}|w|^{2^{*}_{b}}\,dx\right)^{2/2^{*}_{b}}$ $\leq S^{\prime}\left(\int_{B_{1}^{+}}y^{b}|\nabla w|^{2}\,dX+\int_{S_{1}^{+}}y^{b}w^{2}\,d\mathcal{H}^{n}\right),$ (2.1)
$\left(\int_{B_{1}^{+}}y^{b}|w|^{2^{**}_{b}}\,dX\right)^{2/2^{**}_{b}}$ $\leq S\left(\int_{B_{1}^{+}}y^{b}|\nabla w|^{2}\,dX+\int_{B_{1}^{+}}y^{b}w^{2}\,dX\right).$ (2.2)
@@
$r\int_{S_{r}^{+}}y^{b}\left(|\nabla w|^{2}-2\left(\nabla w\cdot\frac{X}{r}\right)^{2}\right)\,d\mathcal{H}^{n}=(n+b-1)\int_{B_{r}^{+}}y^{b}|\nabla w|^{2}\,dX\\ -2\int_{B_{r}^{+}}y^{b}(\nabla w\cdot X)\Delta_{b}w\,dX-2\int_{B_{r}^{\prime}}(\nabla_{x}^{\prime}w\cdot x)\partial_{y}^{b}w\,dx$ (2.3)
@@ 2.2. Regularity of solutions to singular or degenerate elliptic equations
Let $\gamma_{1}\in L^{2}(B_{1}^{+};y^{b})$ and $\gamma_{2}\in L^{2}(B_{1}^{\prime})$ be given, and consider the Neumann problem
$\left\{\begin{array}[]{rll}\mathcal{L}_{b}(w)=&y^{b}\gamma_{1}&\text{ in }B_{R}^{+},\\ \partial_{y}^{b}w=&\gamma_{2}&\text{ on }B_{R}^{\prime}.\end{array}\right$. (2.4)
@@
Let $w\in H^{1}(B_{R}^{+};y^{b})$ be a weak solution to (2.4) with $\gamma_{1}=0$ and $\gamma_{2}\in L^{q}(B_{R}^{\prime})$ with q>n/(1b)q>n/(1-b). Then, for all $r\in(0,R)$ and $\alpha\in(0,1-b-n/q]$ there exists a constant $C=C(n,b,R,r,\alpha)$ such that
$\|w\|_{C^{0,\alpha}}(B_{r}^{+})\leq C\left(\|w\|_{L^{2}(B_{R}^{+};y^{b})}+\|\gamma_{2}\|_{L^{q}(B_{R}^{\prime})}\right).$
Moreover, if $b<0$ and $\gamma_{2}\in L^{q}(B_{R}^{\prime})$ with q>n/bq>-n/b then, for all $r\in(0,R)$ and $\alpha\in(0,-b-n/q]$, there exists a constant $C=C(n,b,r,R,\alpha)$ such that
$\|w\|_{C^{1,\alpha}}(B_{r}^{+})\leq C\left(\|w\|_{L^{2}(B_{R}^{+};y^{b})}+\|\gamma_{2}\|_{L^{q}(B_{R}^{\prime})}\right).$
@@
Let $w\in L^{\infty}(B_{R}^{+})\cap H^{1}(B_{R}^{+};y^{b})$ be a weak solution to (2.4) with $\gamma_{1}=0$ and $\gamma_{2}\in C^{0,\sigma}(B_{R}^{\prime})$ with $\sigma\in(0,1)$. Then, for all $r\in(0,R)$ there exists $\beta\in(0,\sigma)$, $\beta=\beta(n,b,\sigma)$, such that $\partial_{y}^{b}w\in C^{0,\beta}(B_{r}^{+})$. Moreover,
$\|\partial_{y}^{b}w\|_{C^{0,\beta}(B_{r}^{+})}\leq C,$
@@
$\int_{B_{R}}|y|^{b}\nabla w^{*}\cdot\nabla\phi\,dX+\int_{B_{R}}|y|^{b}\gamma_{1}^{*}\phi\,dX=0$ (2.5)
@@
    If $\gamma\in L^{q}(B_{R};|y|^{b})$ with $q>(n+1+b^{+})/2$ and $\alpha\in(0,1)\cap(0,2-(n+1+b^{+})/q]$ then $w\in C^{0,\alpha}(B_{r})$ and there exists a constant $C=C(n,b,R,r,\alpha)$ such that
@@
    If $\gamma\in L^{q}(B_{R};|y|^{b})$ with $q>n+1+b^{+}$ and $\alpha\in(0,1-(n+1+b^{+})/q]$ then $w\in C^{1,\alpha}(B_{r})$ and there exists a constant $C=C(n,b,R,r,\alpha)$ such that
@@ 3. Existence of weak solutions
$\left\{\begin{array}[]{rll}\Delta_{b}^{2}u=&0&\text{ in }B_{1}^{+},\\ \partial_{y}^{b}u=&0&\text{ on }B_{1}^{\prime},\\ \partial_{y}^{b}\Delta_{b}u=&f(\cdot,u)&\text{ on }B_{1}^{\prime},\end{array}\right$. (3.1)
@@
$\int_{B_{1}^{+}}y^{b}\Delta_{b}u\Delta_{b}\phi\,dX=\int_{B_{1}^{\prime}}f(x,u)\phi\,dx$ (3.2)
@@
$\mathcal{H}_{0,S_{1}^{+}}\coloneqq\overline{\left\{C^{\infty}_{c}(\overline{B_{1}^{+}}\setminus\overline{S_{1}^{+}}):\partial_{y}^{b}\phi=0\text{ on }B_{1}^{\prime}\right\}}^{\|\cdot\|_{\mathcal{H}}}.$
@@
$\mathcal{J}(w)\coloneqq\int_{B_{1}^{+}}y^{b}|\Delta_{b}w|^{2}\,dX+\int_{B_{1}^{\prime}}F(x,w)\,dx,$ (3.3)
@@
$\leq\left(\int_{B_{1}^{+}}y^{b}|w_{j}-w_{1}|^{2}\,dX\right)^{1/2}\left(\int_{B_{1}^{+}}y^{b}|\Delta_{b}(w_{j}-w_{1})|^{2}\,dX\right)^{1/2}.$ (3.6)
@@
$0\leq\frac{\mathcal{J}(u+\varepsilon\phi)-\mathcal{J}(u)}{\varepsilon}=\int_{B_{1}^{+}}y^{b}(2\Delta_{b}u\Delta_{b}\phi+\varepsilon|\Delta_{b}\phi|^{2})\,dX+\int_{B_{1}^{\prime}}\frac{F(x,u+\varepsilon\phi)-F(x,u)}{\varepsilon}\,dx.$ (3.7)
@@ 4. Regularity of solutions
For every $\psi\in C_{c}(B_{1}^{+})$ and $R<1$, there exists $\Psi\in\mathcal{H}\cap C^{2,\alpha}(B_{R}^{+})$ such that
@@
As one can readily check, the existence of a weak solution follows from a similar (but simpler) argument to that in the previous section; therefore, we omit the details. The regularity of $\phi$ is a direct consequence of ˜2.10. ∎
@@
Step 1: We begin by considering a smooth cut-off function $\eta\in C^{\infty}_{c}(B_{1})$ such that $\eta\equiv 1$ in $B_{r}$ and $\partial_{y}^{b}\eta=0$, and a function $\Psi\in\mathcal{H}\cap C^{2,\alpha}(B_{R}^{+})$ with $\Psi=0$ on $S_{1}^{+}$, where $r<R<1$ is such that $\operatorname{supp}(\eta)\subset B_{R}$. It is then straightforward to verify that $\eta\Psi\in L^{p}(B_{1}^{\prime})$ and conclude that $\eta\Psi$ is an admissible test function for (3.2) for any such function $\Psi$. Consequently, if we set $w\coloneqq\eta\Delta_{b}u$, we obtain that
$\int_{B_{1}^{+}}y^{b}w\Delta_{b}\Psi\,dX$ $=-\int_{B_{1}^{+}}y^{b}(2\nabla\eta\cdot\nabla\Psi+\Psi\Delta_{b}\eta)\Delta_{b}u\,dX+\int_{B_{1}^{\prime}}f(x,u)\eta\Psi\,dx.$ (4.1)
@@
$\left\{\begin{array}[]{rll}\mathcal{L}_{b}(v)=&0&\text{ in }B_{r}^{+},\\ \partial_{y}^{b}v=&f(\cdot,u)&\text{ on }B_{r}^{\prime}\end{array}\right$. (4.2)
@@
$\int_{B_{r}^{+}}y^{b}\nabla v\cdot\nabla\phi\,dX+\int_{B_{r}^{\prime}}f(x,u)\phi\,dx=0.$ (4.3)
@@
$\|u\|_{L^{q}(B_{r}^{+};y^{b})}+\|u\|_{L^{q}(B_{r}^{\prime})}$ $\leq C\left(\|u\|_{H^{1}(B_{R}^{+};y^{b})}+\|v\|_{H^{1}(B_{R}^{+};y^{b})}\right),$ (4.4)
$\|v\|_{L^{q}(B_{r}^{+};y^{b})}+\|v\|_{L^{q}(B_{r}^{\prime})}$ $\leq C\left(\|u\|_{H^{1}(B_{R}^{+};y^{b})}+\|v\|_{H^{1}(B_{R}^{+};y^{b})}\right).$ (4.5)
@@
$\int_{B_{R}^{+}}y^{b}\nabla u\cdot\nabla\left(\eta_{0,u}^{2}u_{M}^{2^{**}_{b}-2}u\right)\,dX=-\int_{B_{R}^{+}}y^{b}v\left(\eta_{0,u}^{2}u_{M}^{2^{**}_{b}-2}u\right)\,dX.$ (4.8)
Direct computations (see also Lemma 9.1 in [24]) show that, if we set
$K(\tau)\coloneqq\min\left\{\frac{1}{4},\frac{4}{4+\tau}\right\},$
then
$K(2^{**}_{b})\int_{B_{R}^{+}}y^{b}\left|\nabla\left(\eta_{0,u}u_{M}^{(2^{**}_{b}-2)/2}u\right)\right|^{2}\,dX\leq\underbrace{\int_{B_{R}^{+}}y^{b}\nabla u\cdot\nabla\left(\eta_{0,u}^{2}u_{M}^{2^{**}_{b}-2}u\right)\,dX}_{I}\\ +\left(2+K(2^{**}_{b})\frac{2^{**}_{b}+2}{2}\right)\underbrace{\int_{B_{R}^{+}}y^{b}|\nabla\eta_{0,u}|^{2}u_{M}^{2^{**}_{b}-2}u^{2}\,dX}_{II}.$ (4.9)
@@
$\int_{B_{R}^{+}}y^{b}\left|\nabla\left(\eta_{0,u}u_{M}^{(2^{**}_{b}-2)/2}u\right)\right|^{2}\,dX\leq C(n,b,\eta_{0,u})\int_{B_{R}^{+}}y^{b}(|u|^{2^{**}_{b}}+|v|^{2^{**}_{b}})\,dX.$ (4.12)
@@
$p\leq\frac{2^{*}_{b}\cdot 2^{**}_{b}}{2(2+\varepsilon)}+1.$ (4.14)
@@
$\beta_{0}\coloneqq\min\{2^{*}_{b},2^{**}_{b},2+\varepsilon\},$ (4.16)
@@
$\gamma_{0}\coloneqq\frac{2^{*}_{b}\cdot 2^{**}_{b}}{2(p-1)},$ (4.17)
@@
$\|f(\cdot,u)\|_{L^{\gamma_{0}}(B_{r_{0}}^{\prime})}^{\gamma_{0}}\leq C\left(1+\|u\|_{L^{2^{*}_{b}\cdot 2^{**}_{b}/2}(B_{r_{0}}^{\prime})}^{2^{*}_{b}\cdot 2^{**}_{b}/2}\right).$ (4.18)
@@
$K(\beta_{0})\int_{B_{r_{0}}^{+}}y^{b}\left|\nabla\left(\eta_{0,v}v_{M}^{(\beta_{0}-2)/2}v\right)\right|^{2}\,dX\leq-\underbrace{\int_{B_{r_{0}}^{\prime}}\eta_{0,v}^{2}f(x,u)v_{M}^{\beta_{0}-2}v\,dx}_{III}\\ +\left(2+K(\beta_{0})\frac{\beta_{0}+2}{2}\right)\underbrace{\int_{B_{r_{0}}^{+}}y^{b}|\nabla\eta_{0,v}|^{2}v_{M}^{\beta_{0}-2}v^{2}\,dX}_{IV}.$ (4.19)
Similarly to above, we readily obtain that
$|III|$ $\leq\frac{1}{\beta_{0}}\int_{B_{r_{0}}^{\prime}}\eta_{0,v}^{2}|f(x,u)|^{\beta_{0}}\,dx+\frac{\beta_{0}-1}{\beta_{0}}\int_{B_{r_{0}}^{\prime}}\eta_{0,v}^{2}|v|^{\beta_{0}}\,dx,$ (4.20)
$|IV|$ $\leq\int_{B_{r_{0}}^{+}}y^{b}|v|^{\beta_{0}}|\nabla\eta_{0,v}|^{2}\,dX.$ (4.21)

### v3

@@ Abstract.
Corresponding author
Authors ORCID IDs. Donatella Danielli: 0000-0002-8659-6671, Giovanni Gravina: 0000-0001-8985-7964.
@@ 1. Introduction
(1.2) $\Delta_{b}w\coloneqq y^{-b}\operatorname{div}(y^{b}\nabla w)=\frac{b}{y}w_{y}+\Delta w,$
@@
(1.3) $\partial_{y}^{b}w(x,0)\coloneqq\lim_{y\to 0^{+}}y^{b}\partial_{y}w(x,y)$
@@
(1.4) $f(x,u)\coloneqq\lambda_{-}(u^{-})^{p-1}-\lambda_{+}(u^{+})^{p-1},$
@@ 1.1. Motivation and previous works
(1.5) $\left\{\begin{array}[]{rll}\Delta_{b}u=&0&\text{ in }B_{1}^{+},\\ \partial_{y}^{b}u=&\lambda_{+}(u^{+})^{p-1}-\lambda_{-}(u^{-})^{p-1}&\text{ on }B_{1}^{\prime}.\end{array}\right$.
@@
For more details on the (thick) obstacle problem for the bi-Laplacian, we refer the reader to the works of Frehse (see [27] and [28]), Caffarelli and Friedman (see [9]), and Aleksanyan (see [2]). The regularity of solutions to the thin obstacle problem driven by $\Delta^{2}$ was explored by Schild (see [41] and [40]). More recently, versions of the Alt-Caffarelli energy for the bi-harmonic operator were investigated in [19], [18], [37], and [31].
@@ 1.2. Statement of the main results
    If $b\geq 0$, then $(u,v)\in C^{2,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})\times C^{0,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ for all $\alpha\in(0,1-b)$;
@@
    If $b<0$, then $(u,v)\in C^{3,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})\times C^{1,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ for all $\alpha\in(0,-b)$.
@@
    If $b\geq 0$, then $\nabla_{x}^{\prime}v\in C^{0,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ for all $\alpha\in(0,1-b)$;
@@
    If $b<0$, then $\nabla_{x}^{\prime}v\in C^{1,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ for all $\alpha\in(0,-b)$.
@@
Finally, for all $b\in(-1,1)$, there exists $\beta\in(0,1)$ such that $\partial_{y}^{b}v\in C^{0,\beta}(B_{r}^{+}\cup B^{\prime}_{r})$. In particular, if $b=0$ we have that $(u,v)\in C^{3,\beta}(B_{r}^{+}\cup B^{\prime}_{r})\times C^{1,\beta}(B_{r}^{+}\cup B^{\prime}_{r})$ for some $\beta\in(0,1)$.
@@
With Almgren’s frequency formula at our disposal, we follow the pathway outlined in [29]. To be precise, after a preliminary blow-up analysis of the so-called Almgren rescalings of $u$ and $v$ (see ˜6.1), we leverage Weiss- and Monneau-type monotonicity formulas to prove the uniqueness of blow-up solutions. Our findings are summarized in the following theorem, where we use $\mathcal{P}_{\mu}$ to denote a certain class of $\mu$-homogeneous, $b$-harmonic functions (see ˜6.2 below for the precise definition).
@@
for some $\alpha\in(0,1)$ and all $R>0$ as $r\to 0^{+}$. Moreover, $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ cannot both vanish identically.
@@ 2.1. Weighted Sobolev Spaces
This result follows immediately from Lemma 2.5 in [21] once one observes that $|x|<r$ on $B_{r}^{\prime}$.
@@
(2.1) $\left(\int_{B_{1}^{\prime}}|w|^{2^{*}_{b}}\,dx\right)^{2/2^{*}_{b}}$ $\leq S^{\prime}\left(\int_{B_{1}^{+}}y^{b}|\nabla w|^{2}\,dX+\int_{S_{1}^{+}}y^{b}w^{2}\,d\mathcal{H}^{n}\right),$
(2.2) $\left(\int_{B_{1}^{+}}y^{b}|w|^{2^{**}_{b}}\,dX\right)^{2/2^{**}_{b}}$ $\leq S\left(\int_{B_{1}^{+}}y^{b}|\nabla w|^{2}\,dX+\int_{B_{1}^{+}}y^{b}w^{2}\,dX\right).$
@@
(2.3) $r\int_{S_{r}^{+}}y^{b}\left(|\nabla w|^{2}-2\left(\nabla w\cdot\frac{X}{r}\right)^{2}\right)\,d\mathcal{H}^{n}=(n+b-1)\int_{B_{r}^{+}}y^{b}|\nabla w|^{2}\,dX\\ -2\int_{B_{r}^{+}}y^{b}(\nabla w\cdot X)\Delta_{b}w\,dX-2\int_{B_{r}^{\prime}}(\nabla_{x}^{\prime}w\cdot x)\partial_{y}^{b}w\,dx$
@@ 2.2. Regularity of solutions to singular or degenerate elliptic equations
Let $\gamma_{1}\in L^{2}(B_{R}^{+};y^{b})$ and $\gamma_{2}\in L^{2}(B_{R}^{\prime})$ be given, and consider the Neumann problem
(2.4) $\left\{\begin{array}[]{rll}\mathcal{L}_{b}(w)=&y^{b}\gamma_{1}&\text{ in }B_{R}^{+},\\ \partial_{y}^{b}w=&\gamma_{2}&\text{ on }B_{R}^{\prime}.\end{array}\right$.
@@
Let $w\in H^{1}(B_{R}^{+};y^{b})$ be a weak solution to (2.4) with $\gamma_{1}=0$ and $\gamma_{2}\in L^{q}(B_{R}^{\prime})$ with q>n/(1b)q>n/(1-b). Then, for all $r\in(0,R)$ and $\alpha\in(0,1-b-n/q]$ there exists a constant $C=C(n,b,R,r,q,\alpha)$ such that
$\|w\|_{C^{0,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})}\leq C\left(\|w\|_{L^{2}(B_{R}^{+};y^{b})}+\|\gamma_{2}\|_{L^{q}(B_{R}^{\prime})}\right).$
Moreover, if $b<0$ and $\gamma_{2}\in L^{q}(B_{R}^{\prime})$ with q>n/bq>-n/b then, for all $r\in(0,R)$ and $\alpha\in(0,-b-n/q]$, there exists a constant $C=C(n,b,r,R,q,\alpha)$ such that
$\|w\|_{C^{1,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})}\leq C\left(\|w\|_{L^{2}(B_{R}^{+};y^{b})}+\|\gamma_{2}\|_{L^{q}(B_{R}^{\prime})}\right).$
@@
Let $w\in L^{\infty}(B_{R}^{+})\cap H^{1}(B_{R}^{+};y^{b})$ be a weak solution to (2.4) with $\gamma_{1}=0$ and $\gamma_{2}\in C^{0,\sigma}(B_{R}^{\prime})$ with $\sigma\in(0,1)$. Then, for all $r\in(0,R)$ there exists $\beta\in(0,\sigma)$, $\beta=\beta(n,b,\sigma)$, such that $\partial_{y}^{b}w\in C^{0,\beta}(B_{r}^{+}{\cup B^{\prime}_{r}})$. Moreover,
$\|\partial_{y}^{b}w\|_{C^{0,\beta}({B_{r}^{+}\cup B^{\prime}_{r}})}\leq C,$
@@
(2.5) $\int_{B_{R}}|y|^{b}\nabla w^{*}\cdot\nabla\phi\,dX+\int_{B_{R}}|y|^{b}\gamma_{1}^{*}\phi\,dX=0$
@@
    If $\gamma\in L^{q}(B_{R};|y|^{b})$ with $q>(n+1+b^{+})/2$ and $\alpha\in(0,1)\cap(0,2-(n+1+b^{+})/q]$ then $w\in C^{0,\alpha}(B_{r})$ and there exists a constant $C=C(n,b,R,r,q,\alpha)$ such that
@@
    If $\gamma\in L^{q}(B_{R};|y|^{b})$ with $q>n+1+b^{+}$ and $\alpha\in(0,1-(n+1+b^{+})/q]$ then $w\in C^{1,\alpha}(B_{r})$ and there exists a constant $C=C(n,b,R,r,q,\alpha)$ such that
@@ 3. Existence of weak solutions
(3.1) $\left\{\begin{array}[]{rll}\Delta_{b}^{2}u=&0&\text{ in }B_{1}^{+},\\ \partial_{y}^{b}u=&0&\text{ on }B_{1}^{\prime},\\ \partial_{y}^{b}\Delta_{b}u=&f(\cdot,u)&\text{ on }B_{1}^{\prime},\end{array}\right$.
@@
(3.2) $\int_{B_{1}^{+}}y^{b}\Delta_{b}u\Delta_{b}\phi\,dX=\int_{B_{1}^{\prime}}f(x,u)\phi\,dx$
@@
$\mathcal{H}_{0,S_{1}^{+}}\coloneqq\overline{\left\{\phi\in C^{\infty}_{c}(\overline{B_{1}^{+}}\setminus\overline{S_{1}^{+}}):\partial_{y}^{b}\phi=0\text{ on }B_{1}^{\prime}\right\}}^{\|\cdot\|_{\mathcal{H}}}.$
@@
(3.3) $\mathcal{J}(w)\coloneqq\int_{B_{1}^{+}}y^{b}|\Delta_{b}w|^{2}\,dX+\int_{B_{1}^{\prime}}F(x,w)\,dx,$
@@
(3.6) $\leq\left(\int_{B_{1}^{+}}y^{b}|w_{j}-w_{1}|^{2}\,dX\right)^{1/2}\left(\int_{B_{1}^{+}}y^{b}|\Delta_{b}(w_{j}-w_{1})|^{2}\,dX\right)^{1/2}.$
@@
(3.7) $0\leq\frac{\mathcal{J}(u+\varepsilon\phi)-\mathcal{J}(u)}{\varepsilon}=\int_{B_{1}^{+}}y^{b}(2\Delta_{b}u\Delta_{b}\phi+\varepsilon|\Delta_{b}\phi|^{2})\,dX+\int_{B_{1}^{\prime}}\frac{F(x,u+\varepsilon\phi)-F(x,u)}{\varepsilon}\,dx.$
@@ 4. Regularity of solutions
For every $\psi\in C^{\infty}_{c}(B_{1}^{+})$ and $R<1$, there exists $\Psi\in\mathcal{H}\cap C^{2,\alpha}(B_{R}^{+})$ such that
@@
As one can readily check, the existence of a weak solution follows from a similar (but simpler) argument to that in the previous section; therefore, we omit the details. The regularity of $\Psi$ is a direct consequence of ˜2.10. ∎
@@
Step 1: We begin by considering a smooth cut-off function $\eta\in C^{\infty}_{c}(B_{1})$ such that $\eta\equiv 1$ in $B_{r}$ and $\partial_{y}^{b}\eta=0$ on $B_{1}^{\prime}$, and a function $\Psi\in\mathcal{H}\cap C^{2,\alpha}(B_{R}^{+})$ with $\Psi=0$ on $S_{1}^{+}$, where $r<R<1$ is such that $\operatorname{supp}(\eta)\subset B_{R}$. It is then straightforward to verify that $\eta\Psi\in L^{p}(B_{1}^{\prime})$ and conclude that $\eta\Psi$ is an admissible test function for (3.2) for any such function $\Psi$. Consequently, if we set $w\coloneqq\eta\Delta_{b}u$, we obtain that
(4.1) $\int_{B_{1}^{+}}y^{b}w\Delta_{b}\Psi\,dX$ $=-\int_{B_{1}^{+}}y^{b}(2\nabla\eta\cdot\nabla\Psi+\Psi\Delta_{b}\eta)\Delta_{b}u\,dX+\int_{B_{1}^{\prime}}f(x,u)\eta\Psi\,dx.$
@@
(4.2) $\left\{\begin{array}[]{rll}\mathcal{L}_{b}(v)=&0&\text{ in }B_{r}^{+},\\ \partial_{y}^{b}v=&f(\cdot,u)&\text{ on }B_{r}^{\prime}\end{array}\right$.
@@
(4.3) $\int_{B_{r}^{+}}y^{b}\nabla v\cdot\nabla\phi\,dX+\int_{B_{r}^{\prime}}f(x,u)\phi\,dx=0.$
@@
(4.4) $\|u\|_{L^{q}(B_{r}^{+};y^{b})}+\|u\|_{L^{q}(B_{r}^{\prime})}$ $\leq C\left(\|u\|_{H^{1}(B_{R}^{+};y^{b})}+\|v\|_{H^{1}(B_{R}^{+};y^{b})}\right),$
(4.5) $\|v\|_{L^{q}(B_{r}^{+};y^{b})}+\|v\|_{L^{q}(B_{r}^{\prime})}$ $\leq C\left(\|u\|_{H^{1}(B_{R}^{+};y^{b})}+\|v\|_{H^{1}(B_{R}^{+};y^{b})}\right).$
@@
(4.8) $\int_{B_{R}^{+}}y^{b}\nabla u\cdot\nabla\left(\eta_{0,u}^{2}u_{M}^{2^{**}_{b}-2}u\right)\,dX=-\int_{B_{R}^{+}}y^{b}v\left(\eta_{0,u}^{2}u_{M}^{2^{**}_{b}-2}u\right)\,dX.$
Direct computations (see also Lemma 9.1 in [22]) show that there exist constants $C_{1}=C_{1}(2^{**}_{b})$ and $C_{2}=C_{2}(2^{**}_{b})$ such that
(4.9) $\int_{B_{R}^{+}}y^{b}\left|\nabla\left(\eta_{0,u}u_{M}^{(2^{**}_{b}-2)/2}u\right)\right|^{2}\,dX\leq C_{1}\underbrace{\int_{B_{R}^{+}}y^{b}\nabla u\cdot\nabla\left(\eta_{0,u}^{2}u_{M}^{2^{**}_{b}-2}u\right)\,dX}_{I}\\ +C_{2}\underbrace{\int_{B_{R}^{+}}y^{b}|\nabla\eta_{0,u}|^{2}u_{M}^{2^{**}_{b}-2}u^{2}\,dX}_{II}.$
@@
(4.12) $\int_{B_{R}^{+}}y^{b}\left|\nabla\left(\eta_{0,u}u_{M}^{(2^{**}_{b}-2)/2}u\right)\right|^{2}\,dX\leq C(n,b,\eta_{0,u})\int_{B_{R}^{+}}y^{b}(|u|^{2^{**}_{b}}+|v|^{2^{**}_{b}})\,dX.$
@@
(4.14) $p\leq\frac{2^{*}_{b}\cdot 2^{**}_{b}}{2(2+\varepsilon)}+1.$
@@
(4.16) $\beta_{0}\coloneqq\min\{2^{*}_{b},2^{**}_{b},2+\varepsilon\},$
@@
(4.17) $\gamma_{0}\coloneqq\frac{2^{*}_{b}\cdot 2^{**}_{b}}{2(p-1)},$
@@
(4.18) $\|f(\cdot,u)\|_{L^{\gamma_{0}}(B_{r_{0}}^{\prime})}^{\gamma_{0}}\leq C\left(1+\|u\|_{L^{2^{*}_{b}\cdot 2^{**}_{b}/2}(B_{r_{0}}^{\prime})}^{2^{*}_{b}\cdot 2^{**}_{b}/2}\right).$
@@
(4.19) $\int_{B_{r_{0}}^{+}}y^{b}\left|\nabla\left(\eta_{0,v}v_{M}^{(\beta_{0}-2)/2}v\right)\right|^{2}\,dX\leq-C_{1}\underbrace{\int_{B_{r_{0}}^{\prime}}\eta_{0,v}^{2}f(x,u)v_{M}^{\beta_{0}-2}v\,dx}_{III}\\ +C_{2}\underbrace{\int_{B_{r_{0}}^{+}}y^{b}|\nabla\eta_{0,v}|^{2}v_{M}^{\beta_{0}-2}v^{2}\,dX}_{IV},$
where $C_{1}=C_{1}(\beta_{0})$ and $C_{2}=C_{2}(\beta_{0})$ are constants. Similarly to above, we readily obtain that
(4.20) $|III|$ $\leq\frac{1}{\beta_{0}}\int_{B_{r_{0}}^{\prime}}\eta_{0,v}^{2}|f(x,u)|^{\beta_{0}}\,dx+\frac{\beta_{0}-1}{\beta_{0}}\int_{B_{r_{0}}^{\prime}}\eta_{0,v}^{2}|v|^{\beta_{0}}\,dx,$
(4.21) $|IV|$ $\leq\int_{B_{r_{0}}^{+}}y^{b}|v|^{\beta_{0}}|\nabla\eta_{0,v}|^{2}\,dX.$